OpenStudy (anonymous):

@jameskang1015 Solve the quadratic equation by completing the square. =−x2−6x10 First, choose the appropriate form and fill in the blanks with the correct numbers. Then, solve the equation. Round your answer to the nearest hundredth. If there is more than one solution, separate them with commas.

OpenStudy (anonymous):

=-x2-6x10????? thats not it

OpenStudy (anonymous):

x^2-6x-1=0 sorry dumb computer

OpenStudy (anonymous):

its ok one sec

OpenStudy (anonymous):

X=6.16,-0.16 would be it